OpenStudy (anonymous):

10 \[\le\] 4x-2 \[\le\] 22 use interval notation to express the solution set and graph the solution on a set number line>

OpenStudy (anonymous):

[3,6]

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

Add 2 on both sides, then divide by 4. On the left: 10+2=12 and 12/4 = 3. On the right: 22+2=24 and 24/4 = 6.

OpenStudy (anonymous):

thank youuu!!!